How many ounces of vermouth containing 10% alcohol should be added to 20 ounces of gin containing 60% alcohol to make a pitcher of martinis that contains 30% alcohol?
Answer by stanbon(75887)

How many ounces of vermouth containing 10% alcohol should be added to 20 ounces of gin containing 60% alcohol to make a pitcher of martinis that contains 30% alcohol?
------
Equation:
alcohol + alcohol = alcohol
0.10x + 0.60*20 = 0.30(x+20)
--------
Multiply thru by 100:
10x + 60*20 = 30x +30*20
20x = 30*20
x = 30 oz (amt. of 10% vermouth needed)
==========================================
Cheers,
Stan H.
================